# PGD-based optimization of 3D bobsleigh track centerlines from 2D centerlines for simulation applications

### v1

@@ 6 Conclusions and future work
This study proposes a method for generating a 3D bobsleigh track centerline based on a 2D centerline. In accordance with international track design regulations, the method considers factors such as total track length, height, and slope continuity to establish a track generation optimization problem.
Within the selected range of tracks in this study, the proposed algorithm can compute various stylistic centerlines that are consistent with the trend of the actual track centerline, based on either actual or scaled 2D data. The generated centerlines are close to the actual values in terms of length, height difference, and average slope. The maximum errors calculated from actual and scaled 2D data are 1.7%, 3.2%, and 4.1% and 1.1%, 3.5%, and 4.3%, respectively. Moreover, all slope values remain within the specified limits.
By segmenting the track in different ways or adjusting the weight assigned to height difference in the optimization function, multiple stylistically distinct centerlines can be generated. The variety of track centerlines produced by the proposed algorithm can be reasonably utilized to assist in the design of bobsleigh track centerlines.
In the future, based on the wide range of generated national track data, the algorithm proposed in this study can be used to generate stylistically diverse bobsleigh tracks including track sections. Further work can focus on analyzing and constructing representative simulation tracks that incorporate key features from official tracks around the world. This will allow insights obtained from daily training, such as athletes’ driving strategies, to be more universally applicable across different official tracks worldwide. Moreover, exploring and comparing alternative optimization approaches including heuristic, stochastic, and learning-based methods may provide deeper insights into the trade-offs between solution quality, convergence speed, and robustness, relative to the current PGD-based framework.

### v2

@@ 5 Conclusions and future work
This study proposes a method for generating a 3D bobsleigh track centerline based on a 2D centerline. In accordance with international track design regulations, the method considers factors such as total track length, height, and slope continuity to establish a track generation optimization problem. Within the selected range of tracks in this study, the proposed algorithm can compute centerlines whose characteristic parameter trends are consistent with those of the actual track, based on either actual or scaled 2D data. Compared with the actual track, the generated centerlines are close to the actual values in terms of length, height difference, and average slope. Simulation results for all four tracks indicate that the maximum deviation ranges for length, height difference, and average gradient are 0.1%-1.1%, 1.0%-16.4%, and 0.8%-13.5%, respectively.
The sensitivity to segmentation and height-difference weighting differs across tracks, for tracks with real 2D centerline data, higher sensitivity to segmentation and weighting factors is observed, as finer divisions better capture local slope variations, and increasing the weighting factor improves the alignment between the generated and actual elevation profiles. In contrast, tracks based on scaled 2D data exhibit lower sensitivity to both segmentation count and weighting factor, maintaining consistent results under different configurations. Selecting an appropriate number of track segments and weighting factors can effectively reduce deviations between the generated and actual track data, improving the geometric accuracy of the reconstructed centerlines.
In the future, based on the wide range of generated national track data, the algorithm proposed in this study can be used to generate stylistically diverse bobsleigh tracks including track sections. Further work can focus on analyzing and constructing representative simulation tracks that incorporate key features from official tracks around the world. This will allow insights obtained from daily training, such as athletes’ driving strategies, to be more universally applicable across different official tracks worldwide. Moreover, exploring and comparing alternative optimization approaches including heuristic, stochastic, and learning-based methods may provide deeper insights into the trade-offs between solution quality, convergence speed, and robustness, relative to the current algorithm based framework.
@@ 7 Funding
Data availability
The data that support the findings of this study are available from the corresponding author, Jicheng Chen, upon reasonable request.
Appendix A Calculation results
Table 3: Comparison of geometric parameters of the generated and actual Whistler track under different numbers of track segments.
Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
s=15 1286.8 147.3 $4.52\times 10^{-4}$ 0.1703 0.1198
s=26 1286.7 148.2 $3.79\times 10^{-4}$ 0.156 0.1187
s=33 1286.7 148.7 $3.97\times 10^{-4}$ 0.152 0.1182
Actual Track (Whistler) 1278 149.3 $9.76\times 10^{-4}$ 0.204 0.1176
Table 4: Comparison of geometric parameters of the generated and actual Cesana track under different numbers of track segments.
Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
s=10 1356.5 122.2 1.94×10-4 0.1425 0.0939
s=15 1356.4 123.3 1.91×10-4 0.125 0.0926
s=21 1356.7 123.7 2.52×10-4 0.147 0.0920
Actual Track (Cesana) 1359.2 124.5 5.04×10-4 0.183 0.0910
Table 5: Comparison of geometric parameters of the generated and actual Nagano track under different numbers of track segments.
Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
s=10 1702.0 124.5 2.70×10-4 0.1057 0.0734
s=12 1701.5 124.6 1.95×10-4 0.0867 0.0734
s=15 1699.7 126.5 2.04×10-4 0.0878 0.0746
Actual Track (Nagano) 1700 118 - 0.150 0.0700
Table 6: Comparison of geometric parameters of the generated and actual Pyeongchang track under different numbers of track segments.
Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
s=11 1662.0 166.5 2.47×10-4 0.1315 0.1008
s=16 1648.6 170.3 1.87×10-4 0.1170 0.1038
s=22 1677.3 175.0 1.71×10-4 0.1150 0.1049
Actual Track (Pyeongchang) 1659 154 - 0.25 0.0948
Table 7: Comparison of geometric parameters of the generated and actual Whistler track under different height-difference weighting factors.
Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
b=0.5 1287.0 150.8 4.02×10-4 0.1434 0.1167
b=1 1286.7 148.7 3.97×10-4 0.152 0.1182
b=1.5 1286.7 149.1 3.96×10-4 0.157 0.1180
Actual Track (Whistler) 1278 149.3 9.76×10-4 0.204 0.1176
Table 8: Comparison of geometric parameters of the generated and actual Cesana track under different height-difference weighting factors.
Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
b=0.5 1356.3 122.8 1.60×10-4 0.1237 0.0922
b=1 1356.7 123.4 1.62×10-4 0.1252 0.0926
b=1.5 1356.4 123.7 1.63×10-4 0.1265 0.0927
Actual Track (Cesana) 1359.2 124.5 5.04×10-4 0.183 0.0910
Table 9: Comparison of geometric parameters of the generated and actual Nagano track under different height-difference weighting factors.
Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
b=0.5 1702.6 126.4 2.02×10-4 0.0875 0.0744
b=1 1699.7 126.5 2.04×10-4 0.0878 0.0746
b=1.5 1695.4 126.7 2.06×10-4 0.0883 0.0749
Actual Track (Nagano) 1700 118 - 0.150 0.0700
Table 10: Comparison of geometric parameters of the generated and actual Pyeongchang track under different height-difference weighting factors.
Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
b=0.5 1664.5 178.7 1.81×10-4 0.1172 0.1075
b=1 1664.1 177.8 1.81×10-4 0.1172 0.1074
b=1.5 1659.7 177.6 1.82×10-4 0.1174 0.1076
Actual Track (Pyeongchang) 1659 154 - 0.25 0.0948